Using the simplified method, determine the tax-free amount of the following distributions from a qualified pension plan. Contributions, if any, are made with previously taxed dollars.

  1. Person A, age 59, made no contributions to the pension plan and will receive a $500 monthly check for life.

  2. Person B, age 66, made contributions of $23,000 to the pension plan and will receive a monthly check of $1,300 for life.

  3. Person C, age 64, made contributions of $19,000 to the pension plan and will receive monthly payments of $1,200 over her life and the life of her 67-year-old husband.

  4. Person D, age 55, made contributions of $32,000 to the pension plan. He will receive quarterly payments of $5,000 over his life and the life of his 58-year-old wife.
Answer #1
Tax - Free Component Explanation
Person A 0 No contribution made
Person B 23000/210 = $109.52 Anticipated payments is 210
Person C 19000/260 = $73.08 Combined age is 131. Hence, Anticipated payments is 260
Person D 32000/120 = $266.67 Combined age is 113. Hence, Anticipated payments is 120